OpenStudy (anonymous):

A flag pole 30 feet tall casts a shadow of 50 feet on the ground. Find, to the nearest tenth of a degree, the angle of elevation of the sun.

OpenStudy (raden):



OpenStudy (anonymous):

it's been so long since i've done this stuff. retaking a course on it and i covered it more than 2 yearsa go

OpenStudy (raden):

by using defined of tan, we get tan(alpha) = 30/50 = 3/5 = 0.6 solve for alpha by take tan^-1 or arcus(tan(0.6))

OpenStudy (raden):

use ur calculator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got .81

OpenStudy (raden):

i use my caculator, it is 30.9637... degrees

OpenStudy (anonymous):

must've entered it wrong.